shahid sold 3 1/2,7 3/4 and 6 1/2 kg sugar to three customers.how much sugar did he sell in all?
Answer by Cromlix(4381)

3 1/2 = 7/2
7 3/4 = 31/4
6 1/2 = 13/2
7/2 + 31/4 + 13/2
Common Denominator = 4
14 + 31 + 26
------------
4
= 71/4
= 17 3/4kg
Hope this helps.
:-)